The current dividend on a preferred share of Generic Corp. stock is $1. The discount rate is 12%.

If the growth rate of the company is 10%, what would be the price of the stock next year?
Answer #1

Price of the stock is calculated as follows:

Do X (1+9) Price = T-9 Do = current dividend, g = growth rate, r = required rate 1x (1.10) Price = 0.12 - 0.10 Price = 55

So the price is $55